Database last updated on Tue, 03 Feb 2026 13:10:52 +0000 by localhost

Critical warnings from b-files' comments for days between 2022322 and 2022322 (0 sec, 9 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
rl:2022 322 11 18 43 Motor #3: appears stuck with a reference sensor activated.21B32222.254TORON#25405:07:041-1
rl:2022 322 11 41 30 New value set for watermark LOOP.RUN.TIME: 700.000000 millisec.63B32222.254TORON#25405:07:101-1
rl:2022 322 11 41 30 New value set for watermark LOOP.RUN.TIME: 700.000000 millisec.63B32222.254TORON#25405:07:101-1
rl:2022 322 13 43 54 PMT counter failure11B32222.254TORON#25405:07:241-1
da_lo:Brewer failed to respond 5 times31B32222.254TORON#25411:21:541-1
hg:Lamp not on... hg test terminated F1= 0 F4= 1071B32222.254TORON#25414:21:301-1
hg:Lamp not detected f1= 0 f4= 1071B32222.254TORON#25414:21:311-1
hp:Spectrometer alignment trouble.21B32222.254TORON#25417:02:411-1
hg:No intensity detected. Check PMT.11B32222.254TORON#25417:04:061-1

Instrument constant changes for the week before 2022322 (2 sec, 0 rows)

Hg with step corrections larger than 10 steps or multiple corrections larger than 2.5 steps (0 sec, 1 rows)

SerialYYYYJJJHG largemax correction
254202232210815.2

HP with step corrections larger than 20 steps (0 sec, 1 rows)

SerialYYYYJJJHP largemax correction
25420223222263.1

Non-critical warnings from b-files' comments for days between 2022322 and 2022322 (0 sec, 10 rows)

TypeCommentcountpriorityfnamelocationserialtimevalueminmax
rl:2022 322 11 18 42 Reset Tepid -> Tepid from 2022 322 11 18 25 until 2022 322 11 18 37 . Counts: 1,013B32222.254TORON#25405:07:011-1
rl:2022 322 11 18 43 Motor #3: appears stuck with a reference sensor activated.13B32222.254TORON#25405:07:041-1
rl:2022 322 11 18 43 Motor #11: appears stuck with a reference sensor activated.13B32222.254TORON#25405:07:071-1
rl:2022 322 17 03 02 Bad command string. Command string buffer cleared.23B32222.254TORON#25405:07:271-1
rl:2022 322 17 03 02 Bad command string. Command string buffer cleared.23B32222.254TORON#25405:07:271-1
rl:2022 322 17 03 02 <23B32222.254TORON#25405:07:291-1
rl:2022 322 18 01 40 Reset Warm -> Warm from 2022 322 18 01 37 until 2022 322 18 01 38 . Counts: 0,113B32222.254TORON#25405:07:321-1
fi:HOME key pressed499B32222.254TORON#25413:31:491-1
ft:Illegal command (ft) in schedule , ignoring199B32222.254TORON#25416:54:351-1
hg:DEL key pressed299B32222.254TORON#25417:03:591-1

Measurements and tests summary.


_jday_loc_serial_skcdtedhghprssluv
322TORONTO#254inside,menu,Starting schedule inside,Waiting until ZA = 170 now 128.9,Waiting until ZA = 170 now 130.7,Waiting until ZA = 170 now 132.5,Waiting until ZA = 170 now 134.2,Waiting until ZA = 170 now 136,Waiting until ZA = 170 now 137.7,Waiting until ZA = 170 now 139.4,Waiting until ZA = 170 now 141,Waiting until ZA = 170 now 142.6,Waiting until ZA = 170 now 144.2,Waiting until ZA = 170 now 145.7,Waiting until ZA = 170 now 147.1,Waiting until ZA = 170 now 148.5,Waiting until ZA = 170 now 149.8,Waiting until ZA = 170 now 151,Waiting until ZA = 170 now 152.1,Waiting until ZA = 170 now 153,Waiting until ZA = 170 now 153.9,Waiting until ZA = 170 now 154.5,Waiting until ZA = 170 now 155.1,Waiting until ZA = 170 now 155.4,Waiting until ZA = 170 now 155.6,Waiting until ZA = -100 now -101.9825060781
(0 sec)

Number of days since last dome/window cleaning (0 sec, 1 rows)

Missing files (0 sec, 1 rows)

Seriallast datadays ago
25420223230-1

Run/Stop results for last week. (0 sec, 1 rows)

SerialTotalFailed 0Failed darkFailed 1Failed 2Failed 3Failed 4Failed 5Failed double
254803000000

AP results for SB Brewers. (0 sec, 1 rows)

SerialJdayt1 offt1 ont2 offt2 ont3 offt3 onHV offHV on+12V off+12V on+5V off+5V on-12V off-12V on+24V off+24V on+5V sps off+5V sps onSL A offSL A onSL V offSL V onWatts
254202232228.528.429.529.628.728.81389.31388.912.012.05.05.0-12.0-12.024.324.35.05.00.01.60.010.4616.6

Standard lamp intensity analysis (0 sec, 1 rows)

Serialminmaxfilterfixable30-day history
#254196548026281700yes